2. Exercise 5.2

Fred's Hardware and Hobby House expects its sales to increase at a constant rate of 6 percent per year over the next three years. Current sales are $300,000.

Complete the following table by forecasting sales for each of the next three years.

Year

Forecasted Sales

(Dollars)

1   
2   
3   

If sales in 2003 were $180,000 and they grew to $300,000 by 2007 (a four-year period), the actual annual compound growth rate was:

10.76%

18.56%

13.62%

Which of the following are some of the hazards of employing a constant rate of growth forecasting model? Check all that apply.

It assumes that there are no cyclical variations.

It ignores seasonal trends.

It is ill-suited to estimate secular trends.

Forecasted Sales Year (Dollars) $350,000 $282,000 $318,000 sales in 2 80,000 and they gr
Forecasted Sales Year (Dollars) $408,333 f sales in 2 $337,080 80,000 a $265,080
Forecasted Sales Year (Dollars) sales in 2 $476,389 00,000an $357,305 $249,175
Answer #1

F = P *(1+i)^t

Here P = 300000, i =6%, t = 1 to 3

forecasted sale in yr 1 = 300000 * 1.06 = 318000

forecasted sale in yr 2 = 300000 * 1.06^2 = 337080

forecasted sale in yr 3 = 300000 * 1.06^3 = 357304.80 = 357305

If sales in 2003 were 180000 and they grew to 300000 by 2007 (a four-year period)

Let i% be rate of return, then

300000 = 180000*(1+i)^4

i = (300000/180000)^(1/4) - 1 = 1.136219 - 1 = 0.136219 = 13.62%

following are some of the hazards of employing a constant rate

It assumes that there are no cyclical variations.

It ignores seasonal trends.